OpenStudy (anonymous):

EVERYONE GIVE @PHI A MEDAL BECAUSE HE HELPED ME FOR TWO HOURS! HE DESERVES IT!

OpenStudy (tkhunny):

Name the two funds.

OpenStudy (anonymous):

Let x be 5%?

OpenStudy (anonymous):

Let y be 3.5%?

OpenStudy (tkhunny):

Let's work on that just a little. Let x be the amount invested at 5% Let y be the amount invested at 3.5% Do you agree that this sounds more descriptive of exactly what is meant?

OpenStudy (anonymous):

yes. Thank you for helping me by the way and not just telling me the answer.

OpenStudy (anonymous):

so... x+y=413

OpenStudy (tkhunny):

This is why we wrote very clear definitions. x and y are amounts invested. They are NOT interest earned. 413 is interest earned. Thus, x+y CANNOT be 413. They are just not the same thing. We do have a total invested of $10,000. x + y = 10,000 Now, let's talk about interest earned. Do you follow so far?

OpenStudy (anonymous):

x+y=10 000 interest earned is $413..in a year.

OpenStudy (tkhunny):

Okay, how do we use x and y to talk about Simple Interest Earned. We already established that just adding them gives the total amount invested. How shall we use them to give the total interest earned?

OpenStudy (anonymous):

10 000-x+y=413? I'm sorry i'm not really understanding.

OpenStudy (tkhunny):

When in doubt, go read the very clear definitions. Let x be the amount invested at 5% Let y be the amount invested at 3.5% What, then, is the amount earned by fund x? What, then, is the amount earned by fund y?

OpenStudy (anonymous):

x+y=413... wait. out of 10 000 dollars, 5% of it is x. 3.5% of it is y.

OpenStudy (anonymous):

x+y=10 000

OpenStudy (tkhunny):

You did not answer my questions. We already said x+y = 413 is very wrong. Amounts invested are NOT interest earned. Just answer my questions. What is the amount earned by fund x? What is the amount earned by fund y?

OpenStudy (anonymous):

5% for x and 3.5% for y?

OpenStudy (anonymous):

What exactly is interest?

OpenStudy (tkhunny):

No, 0.05 (5%) and 0.035(3½%) are the RATES at which funds are invested. That is not the amount that is earned. You should have this formula, I = Prt, or Interest equals Principal multiplied by the Interest Rate multiplied by the time required. Interest is defined in this way. This requires an understanding of the other three terms. P = the Principal earning interest. In our problem, this is x and y as we have defined them. r = the Interest RATE that is earned or paid. In out problem, these are 0.05 and 0.035 as given in the problem statement. t = the length of time of the investment. The problem statement tells us that this is One Year. It is time to answer the two questions: What is the amount earned (the interest calculated) by fund x? What is the amount earned (the interest calculated) by fund y?

OpenStudy (anonymous):

I have never seen that formula in my life.

OpenStudy (tkhunny):

Well, now you have. One has to wonder why you have been given such a problem if you have no such exposure. Is this a placement exam? If it is a placement exam, you should get this one wrong so that you are placed in the correct class.

OpenStudy (anonymous):

This is just a question from my math textbook. But, I've never seen the formula. Can it not be done without the formula?

OpenStudy (anonymous):

thanks @phi

OpenStudy (tkhunny):

Your question makes no sense. That IS the way to calculate Simple Interest. If you off-handedly decide to discard the only method of calculation as required by the problem statement, you cannot solve the problem or the problem does not actually exist. You can choose which of these dire outcomes is more palatable.

OpenStudy (anonymous):

I'm sorry tkhunny. This question is too hard. I know the first equation, but I'm not able to figure out the second. I understand you are getting annoyed. You can leave if you wish, and maybe @phi can help instead

OpenStudy (phi):

to answer this question, you should watch the first 5 minutes of the video. Can you do that ?

OpenStudy (anonymous):

alright. I'll be back in five minutes. But isn't it wrong to learn something you haven't learned in class yet? Doesn't it misguide you?

OpenStudy (anonymous):

ok. So I watched the first 5 minutes.

OpenStudy (anonymous):

hey @tomhue do you want to help me with this hard question?

OpenStudy (phi):

The main idea is "interest" is money the bank pays you to borrow your money

OpenStudy (anonymous):

so Mengxi had 10 000 dollars to invest. does invest mean loan?

OpenStudy (phi):

Simple interest is how they figure out how much they will pay you. For example, 5% per year simple interest means they will pay you %5 of what you loaned them. Yes, invest means loan in this case

OpenStudy (anonymous):

ok. so Mengxi had a loan of 10 000 from the bank. At the end of the year, she had to pay the bank 10 000 + 413, the interest they asked for.

OpenStudy (phi):

If you invest (or loan) $100 for one year at 5% simple interest rate, you will get back the original 100 plus 5% of 100 (which we calculate at 0.05*100= $5)

OpenStudy (anonymous):

so 10 413. but she also had to give them .05 interest and .35 interest.

OpenStudy (anonymous):

am i on the right track here or no?

OpenStudy (phi):

try not to get confused. Mengxi loans the bank 10,000. The bank pays HER some money to keep that 10,000 for a year. the total they pay her is 413 (plus of course the original 10,000)

OpenStudy (anonymous):

so 10 413

OpenStudy (phi):

yes, but we just keep track of the interest, the 413 Think about this question: if you loaned x dollars to the bank for 1 year at 5% simple interest, how much interest will the bank pay you? any idea ?

OpenStudy (anonymous):

x+.05?

OpenStudy (anonymous):

or i pay the bank that

OpenStudy (anonymous):

Phi, how about we solve an easier question first? I think I'm getting ahead of myself.

OpenStudy (anonymous):

1. Leanne works at a greenhouse. She needs to plant a total of 32 bulbs. two types of bulbs are available. She is asked to plant three times as many crocus bulbs as tulip bulbs. How many of each should she plant.

OpenStudy (anonymous):

Let x be crocus Let y be tulips

OpenStudy (anonymous):

x+y=32

OpenStudy (anonymous):

3x=y

OpenStudy (phi):

you have to be careful about interpreting the English the sentence *** three times as many crocus bulbs as tulip bulbs.*** means you have more crocus bulbs than tulip bulbs in other words, x is bigger than y (so 3x= y is not correct)

OpenStudy (anonymous):

x=32-y 3(32-y)=y 96-3y=y 96=4y 24=y

OpenStudy (anonymous):

what?

OpenStudy (anonymous):

oh so it would be 3y=x?

OpenStudy (phi):

yes x= 3y Unfortunately, the English wording can be confusing... if you have twice as many crocus as tulips, then you c= 2t (or x=2y). same idea if you have three times as many crocus as tulips

OpenStudy (phi):

when I read these problems, I always ask... okay, which is the bigger amount. that helps to figure out is it x=3y or y=3x

OpenStudy (phi):

other than that, you did the problem correctly

OpenStudy (anonymous):

1. x+y=32 x=32-y 2. x=3y 32-y=3y 32=3y+y 32=4y 8=y 1. x+8=32 x=32-8 x=24 There are 24 crocuses and 8 tulips.

OpenStudy (anonymous):

The english language is confusing :)

OpenStudy (anonymous):

I have another question, do you have time?

OpenStudy (phi):

ok

OpenStudy (anonymous):

2. James looks in his TV cabinet an finds some old BETA and VHS tapes. He has 17 tapes in all. He finds that he has three more BETA tapes VHS tapes. How many of each type does he have?

OpenStudy (anonymous):

Let x be # of Beta tapes Let y be # of VHS tapes

OpenStudy (anonymous):

1. x+y=17 y=17-x 2. x=y+3 x=17-x+3 x+x=17+3 2x=20 x=10 1. 10+y=17 y=17-10 y=7 James has 10 beta tapes and 7 vhs tapes

OpenStudy (phi):

perfect

OpenStudy (anonymous):

OH! I FOUND ANOTHER INVESTMENT ONE!

OpenStudy (anonymous):

4. Rehman invests in his summer earnings of $3050. He invests part of the money at 8%/year, and the rest at 7.5%/year. After 1 year, these investments earn $242 in simple interest. how much did he invest at each rate.

OpenStudy (phi):

to do these, you first need to know how to answer this: Let's use $100 as an example. You give the bank $100 you want that $100 back after 1 year. You also want an extra amount of 5% of 100 (the interest) at the end of the year. how much interest will the bank pay you after 1 year?

OpenStudy (anonymous):

so we're giving the bank a loan?

OpenStudy (anonymous):

The bank should give me back 105 dollars.

OpenStudy (phi):

you can think of it that way... but people use different language.. you make a "deposit" or "open an account" but it all means the same thing: the bank has your money (and you don't)

OpenStudy (phi):

The bank should give me back 105 dollars. Yes. How did you figure that out ?

OpenStudy (anonymous):

because .05 multiplied by 100 is 5. So 5 + 100 is 105

OpenStudy (anonymous):

they owe me the initial amount and 5% extra

OpenStudy (phi):

OK. and the amount 0.05*100 is called interest. Next, if you invest (or deposit or loan) the bank x dollars at 5%, how much simple interest do you get after 1 year ?

OpenStudy (anonymous):

x+.05

OpenStudy (phi):

if x were $100, your formula x+0.05 would mean 100+0.05 = 100.05 which does not match the answer of 5. try again.

OpenStudy (phi):

what is *just* the interest for loaning the bank x dollars at 5% for 1year?

OpenStudy (anonymous):

x+5%

OpenStudy (phi):

when we did, what is the interest on $100 at 5% for 1 year, what did you get ?

OpenStudy (anonymous):

OH! x(.05)

OpenStudy (phi):

yes. Now you have the basics for answering the question

OpenStudy (anonymous):

Wow! Thanks! so...I'll try to solve it.

OpenStudy (anonymous):

so.. Let x be 8% Let y be 7.5%

OpenStudy (anonymous):

x+y=242

OpenStudy (anonymous):

wait no. this is wrong

OpenStudy (anonymous):

i already know what x and y

OpenStudy (phi):

the question asks for how much did he invest at each rate. so x and y should be the amounts invested (loaned) at the different rates

OpenStudy (anonymous):

so was my x and y right?

OpenStudy (phi):

*** Let x be 8% Let y be 7.5% *** that is unclear. x is not 8%. x is the amount of money loaned to the bank at 8% interest. at the end of the year the bank will pay interest of 0.08*x

OpenStudy (anonymous):

so x= .08 y= 0.075

OpenStudy (phi):

no. Time to stop and figure it out. x and y are dollars (not percents) x+y is the total dollars given to the bank

OpenStudy (anonymous):

I'm confused again...just when I thought I got it.

OpenStudy (anonymous):

There are 5 numbers: $3050 $242 1 year 7.5% 8%

OpenStudy (phi):

You will get it... review these thoughts: what is *just* the interest for loaning the bank x dollars at 5% for 1year?

OpenStudy (anonymous):

but there are two different interests...

OpenStudy (phi):

You will get it... review these thoughts: what is *just* the interest for loaning the bank x dollars at 5% for 1year?

OpenStudy (anonymous):

he gave the bank 3050

OpenStudy (phi):

yes, but answer the question what is *just* the interest for loaning the bank x dollars at 5% for 1year?

OpenStudy (anonymous):

what do you mean by just?

OpenStudy (anonymous):

x(.05)

OpenStudy (phi):

good. now answer this question what is the interest for loaning the bank x dollars at 8% for 1 year?